BigBite Problems Unsolve, summary what we have done.
Marking point when we have good correlation between the TDC and ADC
..............................................
Monday 3/8/2011 "GOOD" BigBite Data with Beam (run 2598)
Saturday 3/12/2011 at 19:50 "GODD" BigBite Data with cosmics (run 2706). This is the last run with BigBite trigger that we the nice TDC & ADC correlation.
..............................................
Tests that had been done
.............................................
1. Bigbite single Trigger with beam.
result: no self-timming peak. no correlation between TDC & ADC
2. Both TDC from 1190 and F1 have no correlation with ADC
3. Power cycle roc 5 with both immidiate and let it sleep for 5 mins.
result: no change
..............................................
To Check Tomorrow Durring access
...............................................
1, check the VME crate (roc 5) voltage power supply. (suspection that this might gone bad to cause the weird problems)
2. scope out the trigger
3. make another copy of the TDC to the 1877 in roc 8
4. check the proper connections in all possible places.
5. cosmics testing
...............................................